OpenStudy (anonymous):

Palindromes are numbers such as 44 and 383 and 1441 that read the same from right to left as they do from left to right. If a 4-digit number is chosen at random, what is the probability that it will be a palindrome? (a) 1/100 (b) 1/99 (c) 1/90 (d) 1/10 (e) 1/9

OpenStudy (anonymous):

the answer is supposed to be (a) 1/100

OpenStudy (anonymous):

it's a

OpenStudy (anonymous):

but i don't know how

OpenStudy (anonymous):

The first two digits define the palindrome, not counting the first one-digit numbers, we receive \(\frac{1}{90}\)

OpenStudy (anonymous):

(If we consider leading zeroes as non-important)

OpenStudy (anonymous):

Yes, as @LolWolf said it depends on your definition of 4-digit number, which I believe by convention CANNOT start with zeroes...

OpenStudy (anonymous):

In which case it is indeed 1/90

OpenStudy (anonymous):

probability= number of the Palindromes / total number that can be made with 4 digit total number is = 9*10*10*10 number of the Palindromes = 9*10 so probability = 90/9000

OpenStudy (anonymous):

Ah yes right! forgot division! My bad.

OpenStudy (anonymous):

can u explain how u got the numbers that u multiplied please?

OpenStudy (anonymous):

No, wait, you're right, @amir.sat , it is \(\frac{1}{100}\) since the last two digits, which depend on the first two, have a such a probability of equalling the first two.

OpenStudy (anonymous):

for total number u gotta multiply the probability of the numbers that van be in each digit

OpenStudy (anonymous):

and for the Palindromes u gotta multiply on the probability of the first two digit from the left

OpenStudy (phi):

If the answer is supposed to be (a), then they are counting 0120 or 0000 as valid . (Sometimes leading zeros are not allowed) a palindrome will have the form abba where a and b are 0 to 9 once we pick a and b we have picked all 4 digits. the digits ab start with 00,01,...09,10,11,...,20,...99 there are 100 ab pairs. on the other hand there are 0000....9999 or 10,000 4 digit numbers the ratio is 100/10000= 1/100 so 1 out 100 chance of getting a palindrome

OpenStudy (anonymous):

ok thanks